Mothers age was twice the the age of son two years ago.after how many years will she be three times her sons age?
Answer by stanbon(75887)

Mothers age was twice the the age of son two years ago.after how many years will she be three times her sons
------
Equations:
m = 2(s-2)
(m-x) = 3(s+x)
----------------------
Two equations; three variables?
Need more information.
Cheers,
stan H.